If you invest $4,000 at the end of each year at 8% for 40 years, how much will you have in 40 years?
Answer #1

FV =PV*(1+r)^n FV=future value

=4000*(1+0.08)^40   PV=present value

=4000*21.72     r=rate of return

=86898.1     n=time period of investment